SEPARATION OF SPIN SYNCHRONIZED SIGNALS

T. HIGUCHI, G. K. CRAWFORD, R. J. STRANGEWAY AND C. T. RUSSELL

Institute of Geophysics and Planetary Physics, University of California, Los Angeles, CA 90024-1567, U.S.A.

(Received March 16, 1992; revised November 17, 1993)

Abstract. In the measurements of VLF electric fields with the Pioneer Venus spacecraft in sunlight, spin synchronized signals often dominate over the naturally generated emissions. We present a method to separate natural emissions from the several possible sources of noise. Our major objective by this method is not to remove all spin modulation, but to effectively subtract the background noise caused by the identifiable noise sources. Examination of the data shows that the background spin synchronized noise is quite sensitive to θ(n), the angle between the sense axis and the solar direction. We model the observed data as y(n) = w(n)t(n)f(θ(n)) + x(n), where f(θ) represents the phase response of the background noise and x(n) is the estimated natural emissions. t(n) and w(n) are the long-term trend component and time- and phase-independent component of the intensity of the background noise, respectively. The method to decompose y(n) is based on the Bayesian approach which has been recently applied to various inversion problems such as nonstationary time series modeling and image reconstruction. In this procedure, the estimated parameters w(n), t(n), f(θ), and x(n) can be determined automatically. We will describe the Bayesian scheme and its application to the Pioneer Venus VLF electric field data.

Key words and phrases: Time series, Bayesian approach, outlier detection, smoothing, nonlinear modeling.